OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. x − y = 7 y = 3x + 12 A: (2/19, 2/33) B: (-2/19, -33/2) C: (-19/2, -33/2) D: (19/2, 33/2)

OpenStudy (anonymous):

Since you already have one Y solved, plug that to the other equation "x-y=7" and solve for x. Then after you obtained x plug it on the other equation to get y. You will end up having (x,y) point

OpenStudy (anonymous):

ok so... 7=3x+12 then i add seven to both sides right?

OpenStudy (anonymous):

x - (3x+12) = 7 <-- this is how it looks after you plug y into the x - y = 7

OpenStudy (anonymous):

Ok so i know add three to both sides correct?

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

@Mateaus

OpenStudy (anonymous):

first simplify the left side x-(3x+12)

OpenStudy (anonymous):

how would i do that? sorry. im not that good at math

OpenStudy (welshfella):

x - (3x + 12) you need to distribute -1 over the terms in the parentheses = x -1*3x - 1*12 = x - 3x - 12

OpenStudy (anonymous):

first distribute the negative sign before the parenthesis (example: - (2x+3) = -2x-3)

OpenStudy (welshfella):

this = 7 so we have -2x - 12 = 7 solve for x

OpenStudy (anonymous):

19/2

OpenStudy (welshfella):

-2x = 19 x = -19/2

OpenStudy (anonymous):

don't forget to carry on the negative sign :p

OpenStudy (welshfella):

-19/2 as you are dividing 19 by -2

OpenStudy (anonymous):

ok so the awnswer would be (-19/2,-33/2)?

OpenStudy (anonymous):

@Mateaus lol yea

OpenStudy (welshfella):

y = x - 7 = -19/2 - 7 = -19/2 - 14/2 = -33/2

OpenStudy (anonymous):

well since no other x is -19/2 pretty much yes, that's the answer. Although is a good idea to check the other one as well.

OpenStudy (welshfella):

yes - its good for practice to find y

OpenStudy (anonymous):

thank yall both for the help even though i was slow :P medal for @Mateaus and can you medal @welshfella for me please if possable